Gaseous methane reacts with gaseous elemental sulfur to produce gaseous carbon disulfide and hydrogen sulfide gas.

(a) Write the balanced chemical equation. Hint: What is the formula of elemental sulfur?

(b) If 6.20 g of methane and 2.90 g of sulfur are sealed in a 1.00 L vessel, what mass of hydrogen sulfide will be produced if the final temperature is 129.0 °C? Assume 100% yield.

(c) What are the (i) mole fractions, (ii) partial pressures (in bar), and (iii) root-mean-square speed of all gaseous species in the reaction vessel at the end of the reaction (129.0 °C)?




1
Expert's answer
2021-10-04T05:10:30-0400

"2CH_{4(g)} + S_{8(s)} \to 2CS_{2(l)}+4H_2S_{(g)}"


mass of methane = 6.20g

number of moles of methane = 6.20/16 = 0.3875 moles

mass of sulfur = 2.90g

number of moles of sulfur = 2.90/256 = 0.011 moles


limiting reactant is sulfur,

number of moles of h2s = 0.011 × 4 = 0.044 moles

mass of h2s = 0.044 × 34 = 1.55g.